OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 2|3?

OpenStudy (cwrw238):

equation of the circle is x^2 + y^2 = (2/3)^2 x^2 + y^2 = 4/9

OpenStudy (anonymous):

use @cwrw238 equation and plug in a number for x or y and it will give you points.